OpenStudy (anonymous):

The Indianapolis 500 is one of the most exciting events in sports. Each spring, 33 drivers compete in the 500-mile race, sometimes hitting speeds of more than 220 miles an hour. The general formula for computing distance is rate (or speed) multiplied by time, given a steady average speed. The formula may be written d = rt. Answer today's project questions on speed, distance, and time. The driver's average speed to complete all 8 laps is ______ miles per hour. Round to the nearest tenth of a mile per hour.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

1. d = r*t If the race is 8 laps and the total miles is 500, that means that each lap represents 62.5 miles. If the driver does the first 3 laps at 185 miles per hour, we use d = r*t. Each lap is symbolized by d = 62.5 miles. Using the formula, we now have the equation below to find to do the first 3 laps. 3*d/(185) Do the same thing for the next part: 5*d/(200) Add them together and you get t. The answer is 3d/185 + 5d/200 = t ----------------------------- 2. We need to use percentages here. The percent the driver drove 185 miles per hour is 3/8. The percent the driver drove 200 miles per hour is 5/8. 185 * (3/8) + 200*(5/8) = 69.375 + 125 = 194.375 mph To the nearest tenth: 194.4 mph

OpenStudy (anonymous):

This is where Ryan got it I will see if it is correct answer :) http://answers.yahoo.com/question/index?qid=20111018095803AAf0gKU

OpenStudy (anonymous):

thank you!!! is it right?

OpenStudy (anonymous):

yep it is!

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yw! :)

OpenStudy (anonymous):

Bar is this right?